# Couples can be tractable: New algorithms and hardness results for the Hospitals / Residents problem with Couples

### v2

@@ 3 Polynomial-time algorithms
For a connected couple $C_{i}$ of type-a, if $C_{i}$ is not matched in $M$, then by the stability of $M$ and the sub-responsive and sub-complete of the couples, it must be the case all hospital in $A(c_{2i-1})$ are filled with better doctors or all hospitals in $A(c_{2i})$ are filled with better doctors. If $C_{i}$ is connected, then in the former case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the latter we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$.
For a couple $C_{i}$ of type-b, if $C_{i}$ is not matched in $M$, then by the stability of $M$ it must be case that the common hospital $h$ ($\{h\}=A(c_{2i-1})$) is filled with better doctors than $c_{2i-1}$, or every hospital in $A(c_{2i})\setminus\{h\}$ is filled with better doctors than $c_{2i}$ and $h$ has at most a single doctor that is worse than $c_{2i}$ (and also $c_{2i-1}$ as the first case does not hold). In the former case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the latter we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$.
Finally a couple $C_{i}$ of type-c, if $C_{i}$ is not matched in $M$, then by the stability of $M$ it must be the case either all hospital in $A(c_{2i-1})$ are filled with better doctors or all hospitals in $A(c_{2i})$ are filled with better doctors or $h$ ($\{h\}=A(c_{2i-1})\cap A(c_{2i}))$) has a worse doctor than both members, but any other doctor at $h$ is better than both of them and every other acceptable hospital is filled with better doctors than the corresponding member of the couple. In the first case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the second case we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$, while in the third case we add $\{(c_{2i-1},c_{2i})$, $(a_{2i-1},b_{2i-1})$, $(a_{2i},b_{2i})\}$.
@@
Suppose a couple node $c_{k}$ blocks with someone. It cannot block with another $c_{k^{\prime}}$ node, as for each connected couple of type-c, where the edge $(c_{k},c_{k^{\prime}})$ exists and is not in $M^{\prime}$, either one of them is at their best choice (an $a_{i}$ agent), or both of them is at a hospital and in particular at least one of $c_{k}$,$c_{k^{\prime}}$ is at a hospital node that is better than $h^{2}$ ($(\{h\}=A(c_{k})\cap A(c_{k^{\prime}}))$, which by the construction of the preferences is also better than his $c_{k}$ partner. If $c_{k}$ is a member of a half-separable couple, then either $(c_{k},c_{k^{\prime}})\in M^{\prime}$ or the member who can be matched alone is matched to some hospital, which he prefers to $c_{k^{\prime}}$, so it cannot block in this case either.
@@
If $c_{k}$ is matched to his partner’s node $c_{k^{\prime}}$, then we know that the couple is of type-c or is half-separable. Furthermore, in the first case, every hospital node is filled with better partners except maybe $h^{2}$ ($h=H_{k}\cap H_{k^{\prime}})$. If $(c_{k},c_{k^{\prime}})$ is of type-c, then they prefer each other to $h^{2}$, so the edges to $h^{2}$ cannot block. If they are a half-separable couple, then if $c_{k}$ is the member who cannot be assigned alone, then he is with his best partner, so cannot block. If he is the member who can be assigned alone, then each acceptable hospital must be filled with better doctors as $M$ was stable, so $c_{k}$ cannot block with a hospital node again.
If $c_{k}$ is matched to $b_{k}$, by the stability of $M$ and the construction of $M^{\prime}$, we know that each hospital in $H_{k}$ was filled with better doctors in $M$, so $h_{j}^{1},h_{j}^{2}$ is filled with better nodes than $c_{k}$ in $M^{\prime}$ for any $h_{j}\in H_{k}$; or the couple is of type-b and only $h\in H_{k}\cap H_{k^{\prime}}$ has a worse doctor and at most one. Also, in this case $c_{k}$ is the better doctor for $h$ among the couple and so the edge $(c_{k},h^{2})$ does not exists. Hence, the remaining edge $(c_{k},h^{1})$ cannot block, as $h^{1}$ has only better doctors.
@@
Otherwise $c_{k}$ is matched to a hospital node $h_{j}^{1}$/$h_{j}^{2}$. In this case, we claim all nodes corresponding to better hospitals are filled with better doctors, except maybe $c_{k}$’s partner $c_{k^{\prime}}$ might be there and be a worse doctor. Otherwise, if a hospital $h_{j}\notin H_{k}\cap H_{k^{\prime}}$ is not filled with better doctors, then $c_{k}$ could go to $h_{j}$, while his partner remains at the same hospital, which would give a blocking coalition to $M$, contradiction. If there is a $h\in H_{k}\cap H_{k^{\prime}}$ and it is not filled with better doctors, and $h$ is not the worst hospital of $c_{k}$, then the couple is of type-b and $c_{k}$ is better than his partner for $h$. This also implies that we deleted the edge $(c_{k},h^{2})$. As $M$ was stable, $c_{k}$’s partner is the only worse doctor at $h$. Hence, $h^{1}$ is filled with better nodes than $c_{k}$ and the edge $(c_{k},h^{1})$ does not block.
@@
Let $M^{\prime}$ be a stable matching in $I^{\prime}$. We start by showing that for any connected couple $C_{i}$, it holds that $c_{2i-1}$ matched to a hospital node if and only if $c_{2i}$ is too. Suppose $c_{2i-1}$ is matched to some hospital node $h_{j}^{1}$/$h_{j}^{2}$, but $c_{2i}$ is not. Then, as $(c_{2i-1},a_{2i-1})$ does not block, $(a_{2i-1},b_{2i-1})\in M^{\prime}$. As $(b_{2i-1},c_{2i})$ does not block, $(c_{2i},a_{2i})\in M^{\prime}$. But then, $b_{2i}$ is unmatched and $(a_{2i},b_{2i})$ blocks $M^{\prime}$, contradiction. The other case is similar.
If $c_{k}$ is a member of a half-separable couple who cannot be matched alone, then if $c_{k}$ would be matched but his partner $c_{k^{\prime}}$ would not, then the edge $(c_{k},c_{k^{\prime}})$ would block, as $c_{k}$ considers $c_{k^{\prime}}$ as best. So only the designated members of half-separable couples may be matched alone.
It is clear that in $M^{\prime}$, for each hospital node $h_{j}^{1}/h_{j}^{2}$, every doctor at $h_{j}^{1}$ is better for $h_{j}$ than the one (if there is any) at $h_{j}^{2}$. Otherwise that doctor matched to $h_{j}^{2}$ would block with $h_{j}^{1}$, as he prefers $h_{j}^{1}$ to $h_{j}^{2}$.
@@
It is straightforward to see that $M$ is feasible for the hospitals and the single doctors. For the couples, as their preferences are sub-complete, no matter how their two corresponding nodes are matched to hospital nodes, it will correspond to an acceptable assignment of the couple.
@@
Suppose a single doctor $d_{i}$ blocks with a hospital $h_{j}$. But then the edge $(d_{i},h_{j}^{2})$ must block $M^{\prime}$, contradiction.
Suppose a couple $C_{i}$ blocks with two distinct hospitals $h_{j}\neq h_{l}$.
Then, if the couple (or at least one member) is matched to some hospitals already, then it holds that in $I^{\prime}$, either $c_{2i-1}$ or $c_{2i}$ prefer the hospitals $h_{j}^{2}$ or $h_{l}^{2}$ to their assignments, whenever the edges to $h_{j}^{2}$ and $h_{l}^{2}$ exist. If neither of $h_{j},h_{l}$ are a common hospital for the couples, then both edges exists. Hence, $(c_{2i-1},h_{j}^{2})$ or $(c_{2i},h_{l}^{2})$ blocks $M^{\prime}$, contradiction. If $h_{j}=h$ ($\{h\}=A(c_{2i-1})\cap A(c_{2i})$ ), then the edge $(c_{2i-1},h_{j}^{2})$ exists anyway and the same argument works. If, $h_{l}=h$, then the couple is connected, $c_{2i-1}$ only applies to $h$ and is matched there in $M$, so the blocking application would be $(C_{i},h,h)$, contradicting our assumption on $h_{j}\neq h_{l}$. Assume now that the couple $C_{i}$ is unassigned. Then, if neither of them are matched to anywhere, the couple is separable and both $(c_{2i-1},h_{j}^{2})$, $(c_{2i},h_{l}^{2})$ blocks $M^{\prime}$, contradiction. Then, if the edge $(c_{2i},c_{2i-1})$ is in $M^{\prime}$, then as $h_{j}\neq h_{l}$, one of $h_{j}^{2},h_{l}^{2}$ is better for the corresponding member of the couple so the edge $(c_{2i-1},h_{l}^{2})$ or $(c_{2i},h_{l}^{2})$ (both exist) blocks. Otherwise one of the couple nodes must receive the worst partner $b_{2i}$ or $b_{2i-1}$, respectively (If both of them would be matched to their favourite partner $a_{2i-1},a_{2i}$, then $b_{2i}$ would be unmatched so $(a_{2i},b_{2i})$ would block) or remain alone. This couple node, say $c_{2i}$ prefers $h_{l}^{2}$, if it exists, in which case $(c_{2i},h_{l}^{2})$ blocks $M^{\prime}$ contradiction. If the edge to $h_{l}^{2}$ does not exists, then $c_{2i}$ is the better member of the couple for their common hospital $h=h_{l}$ and $C_{i}$ is of type-b. But then, $c_{2i-1}$ only applies to $h$, so $h_{j}=h$, contradicting our assumption on $h_{j}\neq h_{l}$.
Suppose a separable or half-separable couple blocks with a single hospital (so an application $(C_{i},\emptyset,h_{l})$ or $(C_{i},h_{j},\emptyset)$ ). Then, the member who is matched in the blocking coalition either has only edges to hospitals or a $(c_{2i-1},c_{2i})$, which is his worst. So the hospital node $h_{l}^{2}$ or $h_{j}^{2}$ is better for the corresponding node than his partner in $M^{\prime}$, if there is any and as the application blocked $M$, $(c_{2i-1},h_{j})$ or $c_{2i},h_{l})$ blocks $M^{\prime}$, contradiction.
@@
If it of type-c, then as $h$ is the worst for both members, $C_{i}$ must be unassigned. Also, as $C_{i}$ blocks $M$ with $h$, there is at least one worse doctor than $c_{2i-1}$ and a distinct doctor that is worse than $c_{2i}$. So there are at least two doctors that are worse than $c_{2i}$ at $h$, so at least one of them matched to $h^{1}$ in $M^{\prime}$. As $M^{\prime}$ is stable and $(c_{2i-1},h^{2})$ does not block, we get that $c_{2i-1}$ is either matched to $c_{2i}$ or to $a_{2i-1}$ in which case $c_{2i}$ is matched to $b_{2i-1}$. In both cases, we obtain that $(c_{2i},h^{1})$ blocks.
@@
Solvability in polynomial time follows immediately from Theorem 11 as in a Dual Market, each couple is of type-a.
Denote the two classes of hospitals in the Dual Market by $H_{1}$ and $H_{2}$. Denote the set of doctors (both single and member of a couple), who only apply to hospitals in $H_{1}$ by $R_{1}$ and similarly for $R_{2}$.
Furthermore, in this case $A=H_{1}\cup R_{2}$ along with the neighbors $a_{i},b_{i}$ of the couple nodes of $R_{1}$ in their gadget and $B=H_{2}\cup R_{1}$ along with the neighbors $a_{i},b_{i}$ of the couple nodes of $R_{2}$ in their gadget give a bipartition of the nodes, such that neither of them spans any edges, so the constructed instance $I^{\prime}$ is bipartite. Hence, a stable matching always exists and can be found by a Deferred Acceptance algorithm. ∎
@@
We can consider such an instance $I$ as a hrc instance $I^{\prime}$, where each node $v$ with capacity $b(v)$ corresponds to a hospital with capacity $b(v)$ and each edge $e=(u,v)$ corresponds to distinct couples who have a single application $(u,v)$. The preferences of the hospitals are then created according to the corresponding node’s preference lists over the adjacent edges. If there were any loops adjacent to $v$, then we break the tie between the two endpoints of a loop edge $e=(v,v)$ in an arbitrary way (it was one edge, so it was a single entry in $v$’s preference list), and the couple member corresponding to the better end of $e$ will be the better doctor for $v$. It is straightforward to verify that the stable matchings of this hrc instance are exactly the stable matchings in $I$.
@@
In an hrc instance where each couple is of type-a, -b or -c, it holds that in every stable matching, the same set of single doctors are matched and every hospital has the number of doctors. Furthermore, if a hospital is undersubscribed in one stable matching, then it is assigned the same set of doctors in all stable matchings.
@@
First, we note that as the created instance $I^{\prime}$ in Theorem 11 is a sf instance, it is true that each node has the same number of adjacent edges in any stable matching. As the nodes corresponding to single doctors are only adjacent to hospital nodes, and the nodes of the hospitals are only adjacent to doctor nodes, the first part of the theorem immidiately follows.
Let $M,N$ be two stable matchings in $I$ and let $h$ be an underfilled hospital. Take the corresponding stable matchings $M^{\prime},N^{\prime}$ in $I^{\prime}$. In both of them, $h^{2}$ is unmatched, as $h$ was underfilled. They have the same number of adjacent edges for each node, so their symmetric difference $M^{\prime}\triangle N^{\prime}$ is an Euler graph. Hence, it is a union of edge-disjoint cycles. Suppose for the contrary that $h$ has different doctors in $M^{\prime}$ and $N^{\prime}$, so there is a cycle in $M^{\prime}\triangle N^{\prime}$ that is adjacent to $h^{1}$. As none of the edges of the cycle block $M^{\prime}$ or $N^{\prime}$, we get that the preferences must be cyclic, that is each node of the cycle prefers the next node to the previous one in one of the orientations of the cycle. This implies, that there is a node that prefers $h^{1}$ to its partner in $M^{\prime}$ or $N^{\prime}$, say $M^{\prime}$. This node corresponds to a doctor (either single or member of a couple). This doctor is unmatched or at a worse hospital than $h$ in $M$, so by the sub-responsiveness and sub-completeness of the preferences, we get a blocking pair or coalition in any possible case, contradiction. The only case that is not straightforward is the case when this doctor $r=c_{k}$ is a member of a couple that is not assigned in $M$. Then, her partner $c_{k^{\prime}}$ is also not assigned (but is assigned in $N$) and the cycle $C$ must go through the couple’s gadget. As in the cycle $C$, the preferences are cyclic, we get that there is a hospital $h^{\prime}$ that has a worse doctor that $c_{k^{\prime}}$ in $M$. If $h\neq h^{\prime}$, we get that $(c_{k},c_{k^{\prime}},h,h^{\prime})$ blocks $M$, contradiction. If $h=h^{\prime}$, then $h$ has a free place and a worse doctor than $c_{k^{\prime}}$, so $(c_{k},c_{k^{\prime}},h,h)$ blocks, contradiction. ∎

### v3

@@ 3 Polynomial-time algorithms
For a connected couple $C_{i}$ of type-a, if $C_{i}$ is not matched in $M$, then by the stability of $M$ and the sub-responsiveness and sub-completeness of the couples, it must be the case all hospital in $A(c_{2i-1})$ are filled with better doctors or all hospitals in $A(c_{2i})$ are filled with better doctors. If $C_{i}$ is connected, then in the former case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the latter we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$.
For a couple $C_{i}$ of type-b, if $C_{i}$ is not matched in $M$, then by the stability of $M$ it must be case that the common hospital $h$ (where $\{h\}=A(c_{2i-1})$) is filled with better doctors than $c_{2i-1}$, or every hospital in $A(c_{2i})\setminus\{h\}$ is filled with better doctors than $c_{2i}$ and $h$ has at most a single doctor that is worse than $c_{2i}$ (and also $c_{2i-1}$ as the first case does not hold). In the former case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the latter we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$.
Finally a couple $C_{i}$ of type-c, if $C_{i}$ is not matched in $M$, then by the stability of $M$ it must be the case either all hospital in $A(c_{2i-1})$ are filled with better doctors or all hospitals in $A(c_{2i})$ are filled with better doctors or $h$ (where $\{h\}=A(c_{2i-1})\cap A(c_{2i}))$) has a worse doctor than both members, but any other doctor at $h$ is better than both of them and every other acceptable hospital is filled with better doctors than the corresponding member of the couple. In the first case, we add the edges $\{(c_{2i},a_{2i}),(b_{2i},c_{2i-1}),(a_{2i-1},b_{2i-1})\}$ to $M^{\prime}$, in the second case we add $\{(c_{2i-1},a_{2i-1}),(b_{2i-1},c_{2i}),(a_{2i},b_{2i})\}$, while in the third case we add $\{(c_{2i-1},c_{2i})$, $(a_{2i-1},b_{2i-1})$, $(a_{2i},b_{2i})\}$.
@@
Suppose a couple node $c_{k}$ blocks with someone. It cannot block with another $c_{k^{\prime}}$ node, as for each connected couple of type-c, where the edge $(c_{k},c_{k^{\prime}})$ exists and is not in $M^{\prime}$, either one of them is at their best choice (an $a_{i}$ agent), or both of them is at a hospital and in particular at least one of $\{c_{k}$,$c_{k^{\prime}}\}$ is at a hospital node that is better than $h^{2}$ (where $\{h\}=A(c_{k})\cap A(c_{k^{\prime}})$), which by the construction of the preferences is also better than his $c_{k}$ partner. If $c_{k}$ is a member of a half-separable couple, then either $(c_{k},c_{k^{\prime}})\in M^{\prime}$ or the member who can be matched alone is matched to some hospital, which he prefers to $c_{k^{\prime}}$, so it cannot block in this case either.
@@
If $c_{k}$ is matched to his partner’s node $c_{k^{\prime}}$, then we know that the couple is of type-c or is half-separable. Furthermore, in the first case, every hospital node is filled with better partners except maybe $h^{2}$ (where $\{h\}=A(c_{k})\cap A(c_{k^{\prime}}))$. If $(c_{k},c_{k^{\prime}})$ is of type-c, then they prefer each other to $h^{2}$, so the edges to $h^{2}$ cannot block. If they are a half-separable couple, then if $c_{k}$ is the member who cannot be assigned alone, then he is with his best partner, so cannot block. If he is the member who can be assigned alone, then each acceptable hospital must be filled with better doctors as $M$ was stable, so $c_{k}$ cannot block with a hospital node again.
If $c_{k}$ is matched to $b_{k}$, by the stability of $M$ and the construction of $M^{\prime}$, we know that each hospital in $A(c_{k})$ was filled with better doctors in $M$, so $h_{j}^{1},h_{j}^{2}$ is filled with better nodes than $c_{k}$ in $M^{\prime}$ for any $h_{j}\in A(c_{k})$; or the couple is of type-b and only $h$ (where $\{h\}=A(c_{k})\cap A(c_{k^{\prime}})$) has a worse doctor and at most one. Also, in this case $c_{k}$ is the better doctor for $h$ among the couple and so the edge $(c_{k},h^{2})$ does not exists. Hence, the remaining edge $(c_{k},h^{1})$ cannot block, as $h^{1}$ has only better doctors.
@@
Otherwise $c_{k}$ is matched to a hospital node $h_{j}^{1}$/$h_{j}^{2}$. In this case, we claim all nodes corresponding to better hospitals are filled with better doctors, except maybe $c_{k}$’s partner $c_{k^{\prime}}$ might be there and be a worse doctor. Otherwise, if a hospital $h_{j}\notin A(c_{k})\cap A(c_{k^{\prime}})$ is not filled with better doctors, then $c_{k}$ could go to $h_{j}$, while his partner remains at the same hospital, which would give a blocking coalition to $M$, contradiction. If there is a hospital $h$ where $\{h\}=A(c_{k})\cap A(c_{k^{\prime}})$ and $h$ is not filled with better doctors, and $h$ is not the worst hospital of $c_{k}$, then the couple is of type-b and $c_{k}$ is better than his partner for $h$. This also implies that we deleted the edge $(c_{k},h^{2})$. As $M$ was stable, $c_{k}$’s partner is the only worse doctor at $h$. Hence, $h^{1}$ is filled with better nodes than $c_{k}$ and the edge $(c_{k},h^{1})$ does not block.
@@
Let $M^{\prime}$ be a stable matching in $I^{\prime}$. Again, by the worst dummy doctors, we may assume that hospital nodes are full for simplicity (having such a dummy doctor somewhere is equivalent to having that position empty in terms of blocking). We start by showing that for any connected couple $C_{i}$, it holds that $c_{2i-1}$ matched to a hospital node if and only if $c_{2i}$ is too. Suppose $c_{2i-1}$ is matched to some hospital node $h_{j}^{1}$/$h_{j}^{2}$, but $c_{2i}$ is not. Then, as $(c_{2i-1},a_{2i-1})$ does not block, $(a_{2i-1},b_{2i-1})\in M^{\prime}$. As $(b_{2i-1},c_{2i})$ does not block, $(c_{2i},a_{2i})\in M^{\prime}$. But then, $b_{2i}$ is unmatched and $(a_{2i},b_{2i})$ blocks $M^{\prime}$, contradiction. The other case is similar.
If $c_{k}$ is a member of a half-separable couple who cannot be matched alone, then if $c_{k}$ would be matched, but not his partner $c_{k^{\prime}}$, then the edge $(c_{k},c_{k^{\prime}})$ would block, as $c_{k}$ considers $c_{k^{\prime}}$ as best. So only the designated members of half-separable couples may be matched alone.
It is clear that in $M^{\prime}$, for each hospital node $h_{j}^{1}/h_{j}^{2}$, every doctor at $h_{j}^{1}$ is better for $h_{j}$ than the one at $h_{j}^{2}$. Otherwise that doctor matched to $h_{j}^{2}$ would block with $h_{j}^{1}$, as he prefers $h_{j}^{1}$ to $h_{j}^{2}$.
@@
It is straightforward to see that $M$ is feasible for the hospitals and the single doctors. For the couples, we have already shown that no member remains unmatched, who is not allowed to; and as their preferences are sub-complete, no matter how their two corresponding nodes are matched to hospital nodes, it will correspond to an acceptable assignment of the couple.
@@
Suppose a single doctor $d_{i}$ blocks $M$ with a hospital $h_{j}$. But then the edge $(d_{i},h_{j}^{2})$ must block $M^{\prime}$, contradiction.
Suppose a couple $C_{i}$ blocks $M$ with two distinct hospitals $h_{j}\neq h_{l}$.
Then, if the couple (or at least one member) is matched in $M$ to some hospitals already, then it holds that in $I^{\prime}$, either $c_{2i-1}$ or $c_{2i}$ prefer the hospitals $h_{j}^{2}$ or $h_{l}^{2}$ to their assignments, whenever the edges to $h_{j}^{2}$ and $h_{l}^{2}$ exist. If neither of $h_{j},h_{l}$ are a common hospital for the couples, then both edges exists. Hence, $(c_{2i-1},h_{j}^{2})$ or $(c_{2i},h_{l}^{2})$ blocks $M^{\prime}$, contradiction. If $h_{j}=h$ (where $\{h\}=A(c_{2i-1})\cap A(c_{2i})$), then the edges $(c_{2i-1},h_{j}^{2}),(c_{2i},h_{l}^{2})$ exist and the same argument works. If, $h_{l}=h$, then the couple is connected, $c_{2i-1}$ only applies to $h$ and is matched there in $M$, so the blocking application would be $(C_{i},h,h)$, contradicting our assumption on $h_{j}\neq h_{l}$.
Assume now that the couple $C_{i}$ is unassigned in $M$. Then, if neither of them are matched anywhere in $M^{\prime}$, them the couple is separable and both $(c_{2i-1},h_{j}^{2})$, $(c_{2i},h_{l}^{2})$ block $M^{\prime}$, contradiction. If the edge $(c_{2i},c_{2i-1})$ is in $M^{\prime}$, then as $h_{j}\neq h_{l}$, one of $h_{j}^{2},h_{l}^{2}$ is better for the corresponding member of the couple, so the edge $(c_{2i-1},h_{l}^{2})$ or $(c_{2i},h_{l}^{2})$ (both exist) blocks. Otherwise, one of the couple nodes must receive the worst partner $b_{2i}$ or $b_{2i-1}$, respectively (if both of them would be matched to their favourite partner $a_{2i-1},a_{2i}$, then $b_{2i}$ would be unmatched, so $(a_{2i},b_{2i})$ would block), or remain alone. This couple node, say $c_{2i}$ prefers $h_{l}^{2}$, if it exists, in which case $(c_{2i},h_{l}^{2})$ blocks $M^{\prime}$ contradiction. If the edge to $h_{l}^{2}$ does not exists, then $c_{2i}$ is the better member of the couple for their common hospital $h=h_{l}$ and $C_{i}$ is of type-b. But then, $c_{2i-1}$ only applies to $h$, so $h_{j}=h$, contradicting our assumption on $h_{j}\neq h_{l}$.
Suppose a separable or half-separable couple blocks with a single hospital (so an application $(C_{i},\emptyset,h_{l})$ or $(C_{i},h_{j},\emptyset)$ ). Then, the member who is matched in the blocking coalition either has only edges to hospitals or $(c_{2i-1},c_{2i})$, which is his worst. Hence, the hospital node $h_{l}^{2}$ or $h_{j}^{2}$ is better for the corresponding node than his partner in $M^{\prime}$, if there is any and as the application blocks $M$, $(c_{2i-1},h_{j})$ or $(c_{2i},h_{l})$ blocks $M^{\prime}$, contradiction.
@@
If it of type-c, then as $h$ is the worst for both members, $C_{i}$ must be unassigned. Also, as $C_{i}$ blocks $M$ with $h$, there is at least one worse doctor than $c_{2i-1}$ and a distinct doctor that is worse than $c_{2i}$. So there are at least two doctors that are worse than $c_{2i}$ at $h$, so at least one of them matched to $h^{1}$ in $M^{\prime}$. As $M^{\prime}$ is stable and $(c_{2i-1},h^{2})$ does not block, we get that $c_{2i-1}$ is either matched to $c_{2i}$ or to $a_{2i-1}$, in the latter case $c_{2i}$ is matched to $b_{2i-1}$. In both cases, we obtain that $(c_{2i},h^{1})$ blocks.
@@
Solvability in polynomial time follows immediately from Theorem 11 as in a Dual Market, each sub-responsive and sub-complete couple is of type-a.
Denote the two classes of hospitals in the Dual Market by $H_{1}$ and $H_{2}$. For each $i\in\{1,2\}$, denote the set of doctors (both single and a member of a couple), who only apply to hospitals in $H_{i}$ by $R_{i}$.
Furthermore, in relation to the sf instance $I^{\prime}$ constructed in the proof of Theorem 5, in this case $A=H_{1}\cup R_{2}$ along with the neighbors $a_{i},b_{i}$ of the couple nodes of $R_{1}$ in their gadget and $B=H_{2}\cup R_{1}$ along with the neighbors $a_{i},b_{i}$ of the couple nodes of $R_{2}$ in their gadget give a bipartition of the nodes, such that neither of them spans any edges, so the constructed instance $I^{\prime}$ is bipartite. Hence, a stable matching always exists and can be found by a Deferred Acceptance algorithm. ∎
@@
We can consider such an instance $I$ as a hrc instance $I^{\prime}$, where each node $v$ with capacity $b(v)$ corresponds to a hospital with capacity $b(v)$ and each edge $e=(u,v)$ corresponds to distinct connected couples who have a single application $(u,v)$. The preferences of the hospitals are then created according to the corresponding node’s preference lists over the adjacent edges. If there were any loops adjacent to $v$, then we break the tie between the two endpoints of a loop edge $e=(v,v)$ in an arbitrary way (it was one edge, so it was a single entry in $v$’s preference list), and the couple member corresponding to the better end of $e$ will be the better doctor for $v$. It is straightforward to verify that the stable matchings of this hrc instance are exactly the stable matchings in $I$.
@@
In an hrc instance where each couple is of type-a, -b or -c, it holds that in every stable matching,
(i) the same set of single doctors are matched,
(ii) every hospital is assigned the same number of doctors, and
(iii) if a hospital is undersubscribed in one stable matching, then it is assigned the same set of doctors in all stable matchings.
@@
First, we note that by [19, Theorem 2.1] for sf, the sf instance $I^{\prime}$ created in the proof of Theorem 11 satisfies the property that each node has the same number of incident edges assigned to it in any stable matching. As the nodes corresponding to single doctors are only adjacent to hospital nodes, and the nodes of the hospitals are only adjacent to doctor nodes, parts (i) and (ii) of the theorem immediately follow.
For part (iii), let $M_{1},M_{2}$ be two stable matchings in $I$, and let $h$ be a hospital that is undersubscribed in $M_{1}$ (and in $M_{2}$ by part (ii)). Let $M_{1}^{\prime},M_{2}^{\prime}$ be the corresponding stable matchings in $I^{\prime}$, respectively. In both of them, $h^{2}$ is unmatched, as $h$ is undersubscribed in $M_{1},M_{2}$ in $I$. By [19, Theorem 2.1], each of $M_{1}^{\prime},M_{2}^{\prime}$ contains the same number of incident edges at each node, so their symmetric difference $M_{1}^{\prime}\oplus M_{2}^{\prime}$ is an Eulerian graph. Hence, it is a union of edge-disjoint cycles.
Suppose for a contradiction that $h$ is assigned different sets of doctors in $M_{1}$ and $M_{2}$. Then there is a cycle $C$ in $M_{1}^{\prime}\oplus M_{2}^{\prime}$ that is incident to $h^{1}$. As none of the edges of $C$ block $M_{1}^{\prime}$ or $M_{2}^{\prime}$ in $I^{\prime}$, it follows that the preferences in $C$ must be cyclic, that is each node of $C$ prefers the next node to the previous one in one of the orientations of the cycle. This implies that there is a node $v$ of $C$ that prefers $h^{1}$ to its partner in $M_{1}^{\prime}$ or $M_{2}^{\prime}$. Assume by symmetry that $v$ prefers $h^{1}$ to its partner in $M_{1}^{\prime}$. This node $v$ corresponds to a doctor $r$ (either single or a member of a couple in $I$).
If $r$ is single, clearly $(r,h)$ blocks $M_{1}$ in $I$, since $r$ is unmatched in $M_{1}$ or prefers $h$ to $M_{1}(r)$ in $I$, a contradiction. Hence suppose that $r=c_{k}$ is a member of a couple $(c_{k},c_{k^{\prime}})$, without loss of generality. If $c_{k^{\prime}}$ is assigned in $M_{1}$, say to $h^{\prime}$ (possibly $h=h^{\prime}$), then, since $c_{k}$ prefers $h^{1}$ to $M_{1}^{\prime}(c_{k}$), by the sub-responsiveness and sub-completeness of the preferences in $I$, it follows that $(c_{k},c_{k^{\prime}})$ blocks $M_{1}$ with $(h,h^{\prime})$ in $I$, a contradiction.
[…]
Remark 18.
[…]